OpenStudy (anonymous):

Some birds and squirrels were on the branch of a tree. There were 15 animals on the branch with a total of 50 animal feet. How many birds and squirrels were on the tree?

hero (hero):

Algebra problem

hero (hero):

x + y = 15 2x + 4y = 50

OpenStudy (anonymous):

x + y = 15 2x + 4y = 50 solve for x and y (x= birds, y = squrrils)